# Chapter 0 Hadron Production Processes

### v1

@@ 1 Single Pseudoscalar Meson Photoproduction
As mentioned before, kaon physics was the primer for realizing that CP symmetry is violated, with far-reaching implications for elementary particle physics. The existence of two separate kaon decay branches was the first definite signal of competing and interfering processes. Moreover, they enforced the concept of superpositions for elementary particles by the necessity to introduce the $K_{L}$ and $K_{S}$ states of the $K^{0}/\bar{K}^{0}$ system. The kaon database is shown in Fig. 19.
Vector Meson Photoproduction on the Proton
@@
Figure 20: Database for $\gamma p\to\omega p$. Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
There are no vector meson beams, so experiments using modern electromagnetic (EM) facilities attempt to access vector meson nucleon to vector meson nucleon interaction via EM production reactions $ep\to e^{\prime}\text{(vector meson)}~{}p$. Statistics for vector meson photoproduction are very limited. There is a map just for the $\omega$ case (Fig. 20). Data came from JLab, MAMI, ELSA, and GRAAL.
Some vector mesons can, compared to other mesons, be measured with very high precision. This comes from the fact that vector mesons have the same quantum numbers as the photon: $I^{G}(J^{PC})=0^{-}(1^{-~{}-})$. It allows us to apply a Sakurai’s Vector Meson Dominance (VMD) model, assuming that a real photon can fluctuate into a virtual vector meson, which subsequently scatters off a target nucleon (Fig. 21[86].
@@
Figure 21: Schematic diagrams of vector-meson photoproduction (left) and the VMD model (right) in the energy region at threshold experiments. $V$ means vector meson.
@@
High-statistics total cross sections for the vector meson photoproduction at the threshold: $\gamma p\to\omega p$ (from A2 at MAMI [143], ELPH [144], and CBELSA/TAPS) [145], $\gamma p\to\phi p$ (from CLAS [146] and LEPS [147, 148]), and $\gamma p\to J/\psi p$ (from GlueX) [149] allow one to extract the absolute value of the vector meson nucleon scattering length (SL) using the VMD model. The extended analysis of $\Upsilon$-meson photoproduction using quasi-data from the QCD approach is in perfect agreement with the light-meson findings using experimental data [150]
Let us focus on four vector mesons ($\omega$, $\phi$, $J/\psi(1S)$, and $\Upsilon(1S)$) from $q\bar{q}$ nonet, the widths of which are narrow enough to study meson photoproduction at threshold and where data and quasi-data are available (Table 2). To avoid a broad width problem at threshold, one is not considering the $\rho$-meson case to determine vector meson nucleon SL. Furthermore, one will ignore, for example, $\psi^{\prime}(2S)$ due to the difference between the $1S$ and $2S$ states due to “zero” in radial wave functions (WFs). Unfortunately, one cannot go above Quarkonium or $\Upsilon$, whose quark content is $b\bar{b}$. The problem is that actually the Toponium ($T(1S)$), whose quark content is $t\bar{t}$, does not exist. It is due to a large mass of the Theta meson and the $t$-quark decays faster than the quarks form the Theta meson.
@@
Table 2: List of the vector mesons including quark contents and width of them [38]. The four vector mesons, used for the determination of the scattering length as discussed in the text, are marked in blue.
Meson Quark $\Gamma$
Content [MeV]
@@
Figure 22: Comparison of the $|\alpha_{Vp}|$ SLs estimated from threshold vector meson photoproduction on the proton target with VMD model contribution vs the inverse mass of the vector mesons. Input data for phenomenological analyses came from A2 at MAMI (magenta filled circle) [143], ELPH (black filled triangle) [144], and CBELSA/TAPS (black open square) [145] Collaborations for the $\omega$-meson; CLAS (blue filled diamond) [146] and LEPS (green filled square) [147, 148] Collaborations for the $\phi$-meson; and GlueX (red filled star) [149] Collaboration for the $J/\psi$-meson; and quasi-data from Center for Nuclear Femtography (brown open star) [150] for the $\Upsilon$-meson. Analyses results for $\omega$-meson is given at Refs. [143, 144, 151]; for $\phi$-meson is given at Refs. [152, 151]; for $J/\psi$-meson is given at Ref. [153]; and for $\Upsilon$-meson is given at Ref. [154]. The black dashed line is hypothetical following $|\alpha_{Vp}|\propto 1/m_{V}$. $V$ means vector meson.
The “young” vector meson hypothesis may explain the fact that the obtained SL value for the nucleon $\phi$-meson compared to the typical hadron size of approximately $\sim 1~{}\mathrm{fm}$ indicates that the proton is more transparent for the $\phi$-meson compared to the $\omega$-meson and is much less transparent than the $J/\psi$-meson.
@@
In a recent study, the effect of the VMD assumption was studied in the formalism of Dyson-Schwinger equations, which one can consider as an alternative interpretation of the “young age” effect in another (more formal) language [156].
The suggested approach can be employed to evaluate the $J/\psi$-nucleon SLs, replacing the photon by a $J/\psi$-meson (Fig. 21). The results then appear to have the order of several units of $10^{-3}~{}\mathrm{fm}$, $\alpha_{J\psi N}^{(J=1/2)}=(0.2...3.1)\times 10^{-3}~{}\mathrm{fm}$ and $\alpha_{J\psi N}^{(J=3/2)}=(0.2...3.0)\times 10^{-3}~{}\mathrm{fm}$, where $J$ corresponds to the total angular momentum of the $J/\psi$-nucleon system [157].
Future high-quality experiments by EIC and EicC will have the opportunity to evaluate cases for $J/\psi$- and $\Upsilon$-mesons. It allows one to understand the dynamics of $c\bar{c}$ and $b\bar{b}$ production at the threshold. The ability of J-PARC to measure $\pi^{-}p\to\phi n$ and $\pi^{-}p\to J/\psi n$, which are free from the VMD model, is considered.
2 Single Pseudoscalar Meson Electroproduction
An important spectroscopic tool, complementary to photoproduction, is electroproduction of mesons on the nucleon. As indicated in Fig. 23 electron-induced meson production proceeds by the exchange of a virtual photon which allows for the transfer of energy and momentum beyond the on-shell constraints inherent to photoproduction.
Thus, electroproduction data are an important tool to study the properties of non-strange baryons simultaneously in energy and three-momentum transfer as independent variables. Therefore, ongoing PWA fits incorporate the available electroproduction data. The map of $Q^{2}$ dependence of the pion electroproduction data ($\gamma^{\ast}p\to\pi^{0}p$, $\gamma^{\ast}p\to\pi^{+}n$, and $\gamma^{\ast}n\to\pi^{-}p$, no data for $\gamma^{\ast}n\to\pi^{0}n$) before 2009 is shown in Fig. 24. One notes that the CLAS Collaboration produced more than 85% of the world’s pion electroproduction data (this database is still growing), much of which was focused on the mapping of the properties of the $\Delta(1232)$ and higher resonances. Useful comparisons, for instance, will require those involved in this effort to make available all amplitudes obtained in any new determination of ratios $R_{EM}$ and $R_{SM}$ for the transition $N\to\Delta(1232)$ which may be compared with LQCD calculations. Their values are far from those expected of the perturbative regime, $R_{EM}=1$, and $R_{SM}$ is $Q^{2}$ independent [158].
Figure 24: $Q^{2}$ distribution of pion electroproduction data which are now available [159]. Numbers in middle each plot corresponded to number of events. Red filled circles show the number of pion photoproduction events.
3 Strangeness Production on the Nucleon
Strangeness production on the nucleon by excitation of resonances which decay into kaon-hyperon channels is an important spectroscopic tool giving access to the SU(3) flavor structure of baryons. Moreover, such exotic channels like the kaon-hyperon final states are expected to play a key role in identifying hitherto undetected excited states of the nucleon, thus addressing the notorious problem of missing resonances. In [160], pion- and photon-induced $K\Lambda$ reactions were studied within the unitary coupled-channel effective Lagrangian approach. Data on the photoproduction of kaons on the nucleon from the SAPHIR, CLAS and the CBELSA experiment were described by the GiM coupled channels K-matrix approach, also taking into account the full set of all other meson-baryon channels. Thus, a major revision of the complete parameter set was performed. A major goal of those investigations was to address the at that time still open question on the major contributions to the associated strangeness production channels. Since $K\Lambda$ photoproduction data [161, 162] gave an indication for missing resonance contributions, a combined analysis of the $\pi+N\to K\Lambda$ and the $\gamma+N\to K\Lambda$ reactions was expected to identify clearly these states. Assuming small couplings to $\pi N$, these hidden states should not exhibit themselves in the pion-induced reactions and, consequently, in the $\pi N\to K\Lambda$ reaction. The aim of our calculations was to explore to what extent the data available at that time can be explained by known reaction mechanisms without introducing new resonances. Our results for total cross sections are displayed in Fig. 25 and further results on differential cross sections, polarization observables, and angular distributions are found in [160]. As discussed in [160], the SAPHIR [161] and the CLAS [162] data sets, in fact, are leading to two slightly different sets of interaction parameters, reflecting and emphasizing the differences among the two measurements. Below, that point is discussed again.
Figure 25: $\pi^{-}p\to K^{0}\Lambda$ total partial wave cross sections, predicted by parameter set C of Ref. [160], obtained from a fit to the CLAS data [162]. The experimental cross section data are taken from [163, 164, 165].
CLAS-data on $K\Sigma$ production by polarized beams initiated an updated large scale coupled-channels analysis of associated strange production on the nucleon. Based on the effective GiM Lagrangian, a combined CC analysis of $(\pi,\gamma)N\to K\Sigma$ hadro- and photo-production reactions was performed. The analysis covered a center of mass energy range up to $2~{}\mathrm{GeV}$. The central aim was to extract the resonance couplings to the $K\Sigma$ state. Both $s$-channel resonances and $t,~{}u$-channel background contributions are found to be important for an accurate description of angular distributions and polarization observables, assuring a high-quality description of the data. The extracted properties of isospin $I=3/2$ resonances were discussed in detail. In [160], it was found that the $I=1/2$ resonances are largely determined by the non-strangeness channels.
The calculations included 11 isospin $I=1/2$ resonances and 9 isospin $I=3/2$ resonances, respectively. The investigations were extended to the $I=1/2$ and $3/2$ sectors with the parameters fitted to newly published $K\Sigma$ photoproduction data together with the previous $\pi N\to K\Sigma$ measurements in the energy region $\sqrt{s}\leq 2.0~{}\mathrm{GeV}$. The included $K\Sigma$ photoproduction data are those of the $\gamma p\to K^{+}\Sigma^{0}$ published by the LEPS [166, 167, 168], CLAS [169, 170] and GRAAL [171] group, and those of $\gamma p\to K^{0}\Sigma^{+}$ released by the CLAS [172] and CBELSA [173] collaboration, respectively. The SAPHIR data have been left out here because of the known inconsistencies of the $K^{+}\Sigma^{0}$ data [161] with the corresponding CLAS and GRAAL data (for the details, see Ref. [170]). Also, the $K^{0}\Sigma^{+}$ SAPHIR data [161] have much bigger error bars than those of the CBELSA and CLAS groups. The data before 2002 are also no longer used. Results for total cross sections are shown in Fig. 26. Up to a total center-of-mass energy of about $\sqrt{s}=2~{}\mathrm{GeV}$, the data are well described.
Figure 26: Total cross sections for kaon production on the nucleon. Results of the Giessen model [174] are compared to CLAS, CBELSA, and SAPHIR data. Results of the full model calculation are shown in the left panel. Results using only the Born-amplitudes and t-channel meson exchange are displayed in the right panel.
The analysis included all charge channels, $K^{0}\Sigma^{\mp}$ and $K^{\pm}\Sigma^{0}$. A quite satisfactory description was achieved of the $\gamma p\to K^{+}\Sigma^{0}$ data ($\chi^{2}=1.8$) and the $\gamma p\to K^{0}\Sigma^{+}$ data ($\chi^{2}=2.0$). However, the pion-induced strangeness production reactions are described slightly less accurately as indicated by the corresponding $\chi^{2}$ values of $\chi^{2}=$ 4.1, 3.2, and 2.8 for the $\pi^{+}p\to K^{+}\Sigma^{+}$, $\pi^{-}p\to K^{0}\Sigma^{0}$, and $\pi^{-}p\to K^{+}\Sigma^{-}$ reactions, respectively. The parameters have been varied in our fit simultaneously to the $I=1/2$ and $3/2$ sectors. Although the new data are available with reduced total uncertainties, the refitted model parameters were changed only very little. A typical result is displayed in Fig. 27, illustrating the quality of the description on the example of $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The complete set of results, including partial wave cross sections, angular distributions of cross sections, and polarization observables for the full set of $K\Sigma$ exit channels, is found in [174].
Figure 27: Differential cross section of $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The solid (green), dashed (blue) and dotted (magenta) lines are the full model calculation, the model calculation with the $S_{11}(1650)$ and $F_{15}(1680)$ turned off, respectively. The numerical labels denote the center of mass energies in units of GeV. The figure is adapted from Ref. [174].
4 $\eta$-Meson Production
Understanding the dynamics of eta-meson production and, vice versa, the decay of nucleon resonances into the nucleon-eta exit channel is of ongoing interest in hadron spectroscopy. The $\eta$-meson photoproduction on the proton has been measured with high precision by the Crystal Ball Collaboration at MAMI [175]. These high-resolution data provide a new step forward in understanding the reaction dynamics and in the search for a signal from the “weak” resonance states. The main result reported in [175] is a very clean signal for a dip structure around $W=1.68~{}\mathrm{GeV}$, seemingly confirming older data [176, 177, 178, 179]. This raised the question of the origin of that structure, eventually indicating the appearance of a new narrow, possibly exotic, resonance state.
The aim of the study was to extend our previous coupled-channels analysis of the $\gamma p\to\eta p$ reaction by including the data from the new high-precision measurements [175]. The main question is whether the $\eta p$ reaction dynamics can be understood in terms of the established resonance states or whether a new state has to be introduced, thus confirming previous conjectures. A major issue for the analysis is unitarity and a consistent treatment of self-energy effects as visible in the total decay width of resonances. Since the latter are driven by hadronic interactions, the analysis of photo-production data requires the knowledge of the hadronic transition amplitudes as well. Hence, a coupled-channels description as in the Giessen model (GiM) is an indispensable necessity.
As discussed in every detail in Ref. [180] various relevant meson-baryon coupling constants were newly determined at the occasion of this work in large scale coupled-channels calculations. This gave rise to improved constraints on the interaction parameters and the derived resonance parameters, i.e., masses and widths. Representative examples are the mass and width of the $D_{13}(1520)$ resonance, $M=1516\pm 10~{}\mathrm{MeV}$ and $\Gamma=106\pm 4~{}\mathrm{MeV}$, agreeing with and confirming the values obtained earlier by Arndt et al. [181]. It is interesting to note that the mass of this resonance deduced from pion photoproduction tends to be $10~{}\mathrm{MeV}$ lower than the values derived from the pion-induced reactions [182]. The second $D_{13}(1900)$ state has a very large decay width. That state is likely to be related to the $D_{13}(2080)$ two-star resonance, proposed by PDG2024 [38].
The calculations demonstrate a very satisfactory agreement with the experimental data in the whole kinematical region. The first peak is related to the $S_{11}(1535)$ resonance contribution. Similar to the $\pi^{-}p\to\eta n$ reaction, the $S_{11}(1650)$ and $S_{11}(1650)$ states interfere destructively, producing a dip around $W=1.68~{}\mathrm{GeV}$. The coherent sum of all partial waves leads to the more pronounced effect from the dip at forward angles.
In Fig. 29, results for the photon-beam asymmetry $\Sigma$ are compared to the GRAAL data. One can see that even close to the $\eta N$ threshold where the calculations exhibit a dominant $S_{11}$ production mechanism, the beam asymmetry is non-vanishing for angles $\cos(\theta)\geq-0.2$. These results demonstrate that this observable is very sensitive to very small contributions from higher partial waves. At $W=1.68~{}\mathrm{GeV}$ and forward angles the GRAAL measurements show a rapid change of the asymmetry behavior. In [183] this effect was explained by a destructive interference between the $S_{11}(1535)$ and $S_{11}(1650)$ resonances which induces the dip at $W\simeq 1.68~{}\mathrm{GeV}$ in the $S_{11}$ partial wave. Note that the interference between $S_{11}(1535)$ and $S_{11}(1650)$ and the interference between different partial waves are of different nature. The overlapping of the $S_{11}(1535)$ and $S_{11}(1650)$ resonances does not simply mean a coherent sum of two independent contributions, but also includes rescattering (coupled-channel effects). Such interplay is hard to simulate by the simple sum of two Breit-Wigner distributions.
9 Double-Pion Production on the Nucleon

### v2

@@ 4 Strangeness Production on the Nucleon by Kaon Photoproduction
As mentioned before, kaon physics was the primer for realizing that CP symmetry is violated, with far-reaching implications for elementary particle physics. The existence of two separate kaon decay branches was the first definite signal of competing and interfering processes. $K^{0}$ and its antiparticle $\bar{K}^{0}$ are degenerate mass eigenstates which, however, are physically observed in two distinct flavor CP eigenstates, $K^{0}_{1,2}=\frac{1}{\sqrt{2}}(K^{0}\mp\bar{K}^{0})$, which agree up to a CP violating mixing amplitude $\sim 10^{-}3$ with $K_{S,L}$ [27]. This special feature enforced the concept of superpositions for neutral elementary particles by necessitating the introduction of the $K_{L}$ and $K_{S}$ states of the $K^{0}/\bar{K}^{0}$ system. The kaon database is shown in Fig. 22. A similar mixing phenomenon is known for neutral leptons, i.e., neutrinos and antineutrinos.
@@
Figure 22: Database for the Kaon photoproduction data on both proton and neutron targets. Experimental data from the SAID database [112] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data is shown as a stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [154].
Strangeness production on the nucleon through the excitation of resonances that decay into kaon-hyperon channels is an important spectroscopic tool, providing access to the SU(3) flavor structure of baryons. Moreover, such exotic channels, like the kaon-hyperon final states, are expected to play a key role in identifying hitherto undetected excited states of the nucleon, thus addressing the notorious problem of missing resonances. In [169], pion- and photon-induced $K\Lambda$ reactions were studied using the unitary coupled-channel effective Lagrangian approach. Data on the photoproduction of kaons on the nucleon from the SAPHIR, CLAS, and CBELSA experiments were described by the GiM coupled channels K-matrix approach, also taking into account the full set of all other meson-baryon channels. Thus, a major revision of the complete parameter set was performed. A major goal of those investigations was to address the still open question regarding the major contributions to the associated strangeness production channels. Since $K\Lambda$ photoproduction data [170, 171] indicated missing resonance contributions, a combined analysis of the $\pi+N\to K\Lambda$ and the $\gamma+N\to K\Lambda$ reactions was expected to identify these states. Assuming small couplings to $\pi N$, these hidden states should not manifest in the pion-induced reactions and, consequently, in the $\pi N\to K\Lambda$ reaction. The calculations aimed to explore whether the data available at that time could be explained by known reaction mechanisms without introducing new resonances. The results for total cross sections are displayed in Fig. 23 and further results on differential cross sections, polarization observables, and angular distributions are found in [169]. As discussed in [169], the SAPHIR [170] and the CLAS [171] data sets lead to two slightly different sets of interaction parameters, reflecting and emphasizing the differences between the two measurements. Below, that point is discussed again.
Figure 23: $\pi^{-}p\to K^{0}\Lambda$ partial wave cross sections, predicted by parameter set C of Ref. [169], obtained from a fit to the CLAS data [171]. The experimental cross section data are taken from [172, 173, 174]. The figure is adapted from Ref. [169].
CLAS-data on $K\Sigma$ production by polarized beams initiated an updated large scale coupled-channels analysis of associated strange production on the nucleon. Based on the effective GiM Lagrangian, a combined CC analysis of $(\pi,\gamma)N\to K\Sigma$ hadro- and photo-production reactions was performed. The analysis covered a center-of-mass energy range up to $2~\mathrm{GeV}$. The central aim was to extract the resonance couplings to the $K\Sigma$ state. Both $s$-channel resonances and t,ut,~u-channel background contributions are found to be important for an accurate description of angular distributions and polarization observables, ensuring a high-quality description of the data. The extracted properties of isospin $I=3/2$ resonances were discussed in detail. In [169], it was found that the $I=1/2$ resonances are largely determined by the non-strangeness channels.
The calculations included 11 isospin $I=1/2$ resonances and 9 isospin $I=3/2$ resonances, respectively. The investigations were extended to the $I=1/2$ and $3/2$ sectors, with the parameters fitted to newly published $K\Sigma$ photoproduction data, together with the previous $\pi N\to K\Sigma$ measurements in the energy region $\sqrt{s}\leq 2.0~\mathrm{GeV}$. The included $K\Sigma$ photoproduction data are those of the $\gamma p\to K^{+}\Sigma^{0}$ published by the LEPS [175, 176, 177], CLAS [178, 179], and GRAAL [180] groups, as well as those of $\gamma p\to K^{0}\Sigma^{+}$ released by the CLAS [181] and CBELSA [182] collaborations, respectively. The SAPHIR data have been omitted here due to the known inconsistencies of the $K^{+}\Sigma^{0}$ data [170] with the corresponding CLAS and GRAAL data (for details, see Ref. [179]). Also, the $K^{0}\Sigma^{+}$ SAPHIR data [170] have much larger error bars than those of the CBELSA and CLAS groups. The data before 2002 are also no longer used. Results for total cross sections are shown in Fig. 24. Up to a total center-of-mass energy of about $\sqrt{s}=2~\mathrm{GeV}$, the data are well described.
Figure 24: Total cross sections for kaon production on the nucleon. Results of the Giessen model [183] are compared to CLAS, CBELSA, and SAPHIR data. Results of the full model calculation are shown in the left panel. Results using only the Born amplitudes and t-channel meson exchange are displayed in the right panel. The figure is adapted from Ref. [183].
The analysis included all charge channels, $K^{0}\Sigma^{\mp}$ and $K^{\pm}\Sigma^{0}$. A quite satisfactory description was achieved of the $\gamma p\to K^{+}\Sigma^{0}$ data ($\chi^{2}=1.8$) and the $\gamma p\to K^{0}\Sigma^{+}$ data ($\chi^{2}=2.0$). However, the pion-induced strangeness production reactions are described slightly less accurately, as indicated by the corresponding $\chi^{2}$ values of $\chi^{2}=$ 4.1, 3.2, and 2.8 for the $\pi^{+}p\to K^{+}\Sigma^{+}$, $\pi^{-}p\to K^{0}\Sigma^{0}$, and $\pi^{-}p\to K^{+}\Sigma^{-}$ reactions, respectively. The parameters have been varied in the fit simultaneously to the $I=1/2$ and $3/2$ sectors. Although the new data are available with reduced total uncertainties, the refitted model parameters changed only very slightly. A typical result is displayed in Fig. 25, illustrating the quality of the description using the example of the $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The complete set of results, including partial wave cross sections, angular distributions of cross sections, and polarization observables for the full set of $K\Sigma$ exit channels, is found in [183].
Figure 25: Differential cross section of $\pi^{-}p\to K^{0}\Sigma^{0}$ reaction. The solid (green), dashed (blue), and dotted (magenta) lines are the full model calculation, the model calculation with the $S_{11}(1650)$ and $F_{15}(1680)$ turned off, respectively. The numerical labels denote the center of mass energies in units of GeV. The figure is adapted from Ref. [183].
5 Vector Meson Photoproduction on the Proton
For obvious reasons, vector meson beams do not exist, but they can be created in situ by taking advantage of bremsstrahlung, e.g., by scattering a high-energy electron beam off a proton or a nuclear target, as practiced at all major electron facilities. Physically, such experiments require creating $q\bar{q}$ pairs through the interaction of bremsstrahlung photons. A generic Feynman diagram is shown in Fig. 26.
Figure 26: Feynman graph for vector meson photoproduction on a nucleon by bremsstrahlung. A so-called quark transfer process is shown, which is the most likely mechanism for the light vector meson production. The nucleon is denoted as a (qqqqqq) state, the $\omega$ meson as a $q\bar{q}$ state.
Seen from the outside, the process can be explained by Sakata’s VMD hypothesis (illustrated in Fig. 28[97] which in this case applies because massless photons - one of the electro-weak gauge bosons - and light massive vector mesons share not only the same intrinsic quantum numbers $I^{G}(J^{PC})=0^{-}(1^{-~-})$ but also allow the exchange of valence quarks between nucleon and the meson. Thus, at sufficiently high energy, photons may reveal their hadronic content, either in the form of vector mesons, as assumed in VMD, or by coupling to $q\bar{q}$ pairs in a vector meson configuration. The hit proton is in an excited state, which subsequently decays by $\omega$ emission. Examples of $\omega$-production in bremsstrahlung experiments at CLAS@JLab and MAMI@Mainz are found in [184] and in [185], respectively. The JLab measurement was performed with a bremsstrahlung photon beam produced by a continuous electron beam of energy 4.1+ GeV hitting a gold foil, combined with a bremsstrahlung tagging system. $\omega$ production was identified from the $\gamma p\to p+X$ missing mass spectrum. For the Mainz experiment, the energy-tagged bremsstrahlung-photon beam produced by the MAMI electron beam was used, with the Crystal Ball (CB) serving as a central spectrometer, and TAPS installed as a forward spectrometer for measuring the chain of processes $\gamma p\to\omega p\to\pi^{0}\gamma p\to 3\gamma p$. The CBELSA/TAPS experiment at Bonn uses similar techniques, see, e.g., [186].
Figure 27: Database for $\gamma p\to\omega p$. Experimental data from the SAID database [112] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data is shown as a stacked histogram for cross sections data (Light shaded ) and polarization data (dark shaded). The figure is adapted from Ref. [154].
In Fig. 27, the data available from JLab, MAMI, ELSA, and GRAAL for the $\omega$ case are shown.
@@ 5 Vector Meson Photoproduction on the Proton
Figure 28: Schematic diagrams of vector-meson photoproduction (left) and the VMD model (right) in the energy region of threshold experiments. $V$ means vector meson.
@@
High-statistics total cross sections for vector meson photoproduction at the threshold: $\gamma p\to\omega p$ (from A2 at MAMI [185], ELPH [187], and CBELSA/TAPS) [188], $\gamma p\to\phi p$ (from CLAS [189] and LEPS [190, 191]), and $\gamma p\to J/\psi p$ (from GlueX) [192] allow us to constrain the modulus of the vector meson nucleon scattering length (SL) by using the VMD model to extrapolate to the on-shell point. The extended analysis of $\Upsilon$-meson photoproduction using QCD quasi-data is in perfect agreement with the light-meson findings using experimental data [193]. Interestingly, similar to the high-energy limit, the near-threshold production amplitude was found theoretically to be factorizable into gluonic generalized parton distributions (GPD) and a quarkonium distribution amplitude.
Let us focus on four vector mesons ($\omega$, $\phi$, $J/\psi(1S)$, and $\Upsilon(1S)$) from the $q\bar{q}$ nonet, the widths of which are narrow enough to study meson photoproduction at threshold, and for which data and quasi-data are available (Table 2). To avoid a broad width problem at the threshold, the $\rho$-meson is not considered a meaningful candidate for the determination of the vector meson nucleon SL. Furthermore, $\psi^{\prime}(2S)$ is not considered because of the disturbing effects from the additional node in the $2S$ radial wave function (WFs). Until recently, it was expected that investigations could not be extended beyond the $\Upsilon$ $b\bar{b}$ region because the $t\bar{t}$ state $T(1S)$ seemed not to exist. The situation changed with recent CMS observations of a near-threshold enhancement in top quark pair production [194].
@@
Table 2: List of the vector mesons including quark contents and their widths [41]. The four vector mesons, used for the determination of the scattering length as discussed in the text, are marked in blue.
Vector Quark $\Gamma$
Meson Content [MeV]
@@
Figure 29: Comparison of the $|\alpha_{Vp}|$ SLs estimated from threshold vector meson photoproduction on the proton target with VMD model contribution vs the inverse mass of the vector mesons. Input data for phenomenological analyses came from A2 at MAMI (magenta filled circle) [185], ELPH (black filled triangle) [187], and CBELSA/TAPS (black open square) [188] Collaborations for the $\omega$-meson; CLAS (blue filled diamond) [189] and LEPS (green filled square) [190, 191] Collaborations for the $\phi$-meson; and GlueX (red filled star) [192] Collaboration for the $J/\psi$-meson; and quasi-data from Center for Nuclear Femtography (brown open star) [193] for the $\Upsilon$-meson. Analyses results for $\omega$-meson is given at Refs. [185, 187, 195]; for the $\phi$-meson is given at Refs. [196, 195]; for $J/\psi$-meson is given at Ref. [197]; and for $\Upsilon$-meson is given at Ref. [198]. The black dashed line is hypothetical following $|\alpha_{Vp}|\propto 1/m_{V}$. $V$ means vector meson.
The suggested approach can be applied to evaluate the $J/\psi$-nucleon SLs, replacing the photon with a $J/\psi$-meson (Fig. 28). The results then appear to have the order of several units of $10^{-3}~\mathrm{fm}$, $\alpha_{J\psi N}^{(J=1/2)}=(0.2...3.1)\times 10^{-3}~\mathrm{fm}$ and $\alpha_{J\psi N}^{(J=3/2)}=(0.2...3.0)\times 10^{-3}~\mathrm{fm}$, where $J$ corresponds to the total angular momentum of the $J/\psi$-nucleon system [199].
Future high-quality experiments by EIC and EicC will have the opportunity to evaluate cases for $J/\psi$- and $\Upsilon$-mesons. It allows one to understand the dynamics of $c\bar{c}$ and $b\bar{b}$ production at the threshold. The ability of J-PARC to measure $\pi^{-}p\to\phi n$ and $\pi^{-}p\to J/\psi n$, independent of the VMD model, is considered. A new flavor is introduced into these investigations by lattice QCD studies of the HAL-QCD collaboration. In [200], low-energy interactions $N-J/\psi$ and $N-\eta_{c}$ are based on (2+1) flavor configurations with a nearly physical pion mass, $m_{\pi}=146$ MeV. The interactions are found to be attractive in all distances and possess a characteristic long-range tail consistent with the two-pion exchange potential.
6 Electroproduction of Single Pseudoscalar Mesons
An important spectroscopic tool, complementary to photoproduction, is the electroproduction of mesons on the nucleon. As indicated in Fig. 30, electron-induced meson production proceeds by the exchange of a virtual photon, which allows for the transfer of energy and momentum beyond the on-shell constraints inherent to photoproduction.
Thus, electroproduction data are an important tool for studying the properties of non-strange baryons simultaneously in energy and three-momentum transfer as independent variables. Therefore, ongoing PWA fits incorporate the available electroproduction data. The map of $Q^{2}$ dependence of the pion electroproduction data ($\gamma^{\ast}p\to\pi^{0}p$, $\gamma^{\ast}p\to\pi^{+}n$, and $\gamma^{\ast}n\to\pi^{-}p$, no data for $\gamma^{\ast}n\to\pi^{0}n$) before 2009 is shown in Fig. 31. One notes that the CLAS Collaboration produced more than 85% of the world’s pion electroproduction data (this database is still growing), much of which was focused on mapping the properties of the $\Delta(1232)$ and higher resonances. Useful comparisons, for instance, will require those involved in this effort to make available all amplitudes obtained in any new determination of ratios $R_{EM}$ and $R_{SM}$ for the transition $N\to\Delta(1232)$, which may be compared with LQCD calculations. Their values are far from those expected in the perturbative regime, $R_{EM}=1$, and $R_{SM}$ are $Q^{2}$ independent [201].
Figure 31: $Q^{2}$ distribution of pion electroproduction data, which are now available [202]. Numbers in the middle of each plot corresponded to the number of events. Red-filled circles show the number of pion photoproduction events.
9 Double-Pion Production on the Nucleon
1 Overview on Double-Pion Production by Pion Beams
Early studies of the dynamics of the $N\pi\pi$ final state were based on bubble chamber data on the reaction $\pi p\to\pi\pi N$ collected in non-polarized experiments by the Berkeley, Saclay, and Rutherford laboratories (Fig. 32). 241,214 Bubble Chamber events have been analyzed in Isobar-model PWA at $W=1320-1930~\mathrm{MeV}$ by the Virginia Tech group led by Richard Arndt [203]. A summary of the number of Bubble Chamber events is given in Fig. 32.
Partial-wave inelasticities for elastic pion-nucleon scattering were determined with the aid of experimental data on $\pi N\to\pi\pi N$ processes in the beam-momentum range $300~\mathrm{MeV/c}<\mathrm{P_{beam}}<500~\mathrm{MeV/c}$ [204]. Respective partial wave results are displayed in Fig. 32.
However, an obstacle for a much-desired joint PWA analysis is the inconsistencies that, at present, inhibit merging $\pi N\to\pi N$ and $\pi N\to\pi\pi N$ databases. Problems exist specifically for small inelasticities (Fig. 32) While the largest inelastic cross section related to the $P_{11}$ Roper resonance is in excellent agreement, it is well described in SAID-SP06 for $\pi N\to\pi N$ [166], see Fig. 32.
Besides photo-production and electro-production on the nucleon [203, 166, 202, 205] double-pion production was also measured systematically in NN scattering at WASA@CELSIUS and WASA@COSY [206, 207, 89].
@@ 1 Overview on Double-Pion Production by Pion Beams
Present coupled-channel models for double-meson production involve approximations, e.g., describing three-body interactions by effective two-body configurations. The caveats are well understood on a formal level, but their numerical implementation is a highly non-trivial task. Improvements and optimization strategies are ongoing tasks in coupled-channel research, as is a deeper look into the developments documented in the discussions of competing approaches, e.g.[209, 208, 210, 117, 119, 102, 211]. A practical approach is discussed in the following paragraph.
2 Double-Pion Production in Pion-Proton Reactions